OpenStudy (anonymous):

Two pizzas and a salad cost $29.00. Three pizzas and 2 salads cost $46.00. How much is a pizza? How much is a salad? Write the system of equations that would be used to solve this problem.

OpenStudy (anonymous):

2x + y = 29 3x +2y = 46 y = 29 - 2x 3x + 2(29-2x) = 46 -x = -12 x = 12 y = 5

OpenStudy (anonymous):

the substitute the value of x to the equation to get the y.